Ordered: Tobramycin sulfate 60mg tid IV for a 40-lb child. Safe doses range: 6 to 7.5 mg/kg/day in 3 divided doses.

Is the order safe?
Answer #1

ANSWER.

* Calculation of safe dose for this child.

          Ordered dose = 60 mg tid.

tid is three times a day

    so

         Ordered dose = 60 * 3

                                = 180 mg per day.

          Weight of the patient = 40 lb.

Convert lb to kg for calculation

             1 lb = 0.454 kg

       SO

            Weight of the child = 40 * 0.454

                                           = 18.2 kg.

   Safe recommended dose range = 6 to 7.5 mg/kg/day.

       Safe dose for this child = 6 * 18.2 to 7.5 * 18.2

                                             = 109.2 mg to 136.5 mg per day.

    * So the ordered dose (180 mg/day) is not safe for this child.